OpenStudy (helpblahblahblah):

ILL GIVE MEDALS: what percent of 65 is 13?

OpenStudy (lucaz):

x% of 65 = 13 (x/100)*65 = 13 solve for x

OpenStudy (anonymous):

The formula to your percentage problem is percent over 100 equals is over of. Meaning that is, is considered as a numerator and of is considered as a denominator. So, you place 13 as the numerator and 65 as the denominator. You multiply 13 by 100 and then you divide the answer by 65. --- * ---- 13 * 100 / 65 = 20. 20 is your answer.

OpenStudy (anonymous):

20

OpenStudy (mertsj):

Ok, blah blah blah, you have been shown how to do these problems at least 3 times now. Time for you to do one on your own. Post your answer next time and we will check it for you.

OpenStudy (helpblahblahblah):

ohh thanks @britneyspears and @Mertsj i just dont get it

OpenStudy (anonymous):

You're welcome!! Xo!

OpenStudy (mertsj):

Well then, it does absolutely no good to have someone else do your work for you. Do you think you will get it if you don't do the work?

OpenStudy (helpblahblahblah):

no

OpenStudy (mertsj):

So you should try the next one. If you don't get it post your work and we can guide you. But you need to at least start the problem.

OpenStudy (helpblahblahblah):

okay sorry

OpenStudy (mertsj):

Don't be sorry. Just remember. You are the ONLY person who can get an education for yourself.

OpenStudy (helpblahblahblah):

okay